# MathOPEval: A Fine-grained Evaluation Benchmark for Visual Operations of MLLMs in Mathematical Reasoning

### v1

@@ 5.3 In-Depth Analysis
Points vs. Lines: Regarding the generation setting for multi-modal code annotations in geometric problems, we further categorize the annotations into two types: angle markings and line labels. As illustrated in Figure 4, angle annotations demonstrate notably higher scores compared to line annotations. We attribute this performance difference to the varying complexity of placement requirements: angle markings only need to be positioned near the identified vertex, whereas line labels require to be placed at the midpoint of the line segment. This more stringent positioning requirement for line labels contributes to the increased difficulty in their accurate placement.
5.4 Error Analysis
@@
We conduct a manual analysis on 100 randomly selected outputs containing reasoning processes, equally divided between 50 correct and 50 incorrect processes. Our examination reveals concerning patterns in the models’ performance: among the correct answers, only 33% demonstrates both accurate results and valid reasoning processes, while 17% achieves correct answers through wrong reasoning - indicating that correct outputs do not necessarily reflect sound problem-solving capabilities. We also identify four distinct categories of failures, with visual perception errors emerging as the predominant issue, accounting for 86% of all errors. This remarkably high percentage highlights a crucial limitation in the models’ ability to accurately process and interpret visual elements. Both instruction comprehension errors and output format violations each contribute to 6% of the failures, while reasoning process errors constitute the remaining 2%. This distribution of errors suggests that visual perception mechanisms require substantial improvement for reliable visual problem-solving.
6 related work
@@
Multimodal Mathematical Assessment: With the rapid development of MLLMs, numerous benchmarks have emerged to evaluate visual mathematical reasoning capabilities(Wang et al., 2024; He et al., 2024; Yue et al., 2024; Zhang et al., 2024a; Wang et al., 2024; Sun et al., 2024; Qiao et al., 2024; Zhou et al., 2024; Xu et al., 2023; Masry et al., 2022). Previous work (Chen et al., 2022; Seo et al., 2015; Cao & Xiao, 2022), such as Inter-gps (Lu et al., 2021) and GeoQA (Chen et al., 2021), focuses on geometric problems, while others (Lu et al., 2024; Zhang et al., 2024b) cover broader tasks including function graphs and statistical charts. Though recent work (Hu et al., 2024a; Cheng et al., 2024; Lee et al., 2024) introduces a “multi-modal input-multi-modal output” paradigm, existing evaluations still assess only final answer accuracy, overlooking the quality of intermediate visual operations during reasoning.
@@
Visual Programming: The decomposition of complex visual tasks into manageable steps through programming interfaces has demonstrated significant potential in advancing visual reasoning capabilities(Yao et al., 2023; Yang et al., 2023; Zeng et al., 2023; Hu et al., 2024b). The core idea behind this approach is to leverage the logical and precise nature of computer programs to transform complex visual problems, which traditionally rely on human intuition and experience, into a series of clear, step-by-step operations. Among them, Visprog (Gupta & Kembhavi, 2023) and ViperGPT (Surís et al., 2023) showcase the effectiveness of using MLLMs to generate Python code for sequential visual operations. SKETCHPAD (Hu et al., 2024a) further enables dynamic visual manipulations based on intermediate results, achieving particular success in geometric reasoning through visual operations like drawing auxiliary lines via code. This progress inspires our work to move beyond evaluating just final mathematical solutions and toward assessing the quality and appropriateness of intermediate visual operations in the reasoning process.
7 conclusion and Future Work
@@
In this paper, we presented the first fine-grained evaluation benchmark for visual operations in mathematical reasoning tasks. Our framework comprehensively assesses four fundamental visual operations across five types of mathematical visualizations, providing a systematic approach to evaluating MLLMs’ capabilities. Through extensive experiments, we identified significant performance gaps between current models and human, particularly in function plots and multi-modal code modification task. Additionally, we observed that while in-context learning can enhance performance, its effectiveness varies across different tasks and requires careful consideration of shot count. We hope our framework and findings will facilitate the development of more capable models in visual mathematical reasoning.
This work initiates the evaluation of visual operations of MLLMs in mathematical reasoning tasks, paving the way for future research. Several directions remain open: (1) expanding the evaluation scope to cover more multimodal reasoning skills such as planning, reflection, and error correction in mathematical problems; and (2) exploring how this benchmark can be leveraged to align MLLMs and enhance their multimodal reasoning capabilities in mathematical scenarios.
Appendix A Prompt Design
For different evaluation scenarios, we design specific prompts to guide the model’s response. We consider two main question types: multiple-choice questions and free-form generation tasks.
For multiple-choice questions, the prompts are structured as follows:
  • Direct:
    Question: {question}
    Choices:
    {formatted_choices}
    Please directly output the correct option.
  • CoT:
    Question: {question}
    Choices:
    {formatted_choices}
    Let’s think step by step. Please equip the correct
[…]
  • DCoT:
    Question: {question}
    Choices:
    {formatted_choices}
    Describe the image information relevant to the
    question.
[…]
    end of your response.
  • VCoT:
    Question: {question}
    Choices:
    {formatted_choices}
    Visualize the state after each reasoning step.
[…]
    end of your response.
For free-form generation tasks, the prompts follow similar patterns:
  • Direct:
    Question: {question}
    Please directly output your code using markdown
    format.
  • CoT:
    Question: {question}
    Let’s think step by step. Please wrap your code
    using markdown format at the end of your response.
  • DCoT:
    Question: {question}
    Describe the image information relevant to the
    question.
    Please wrap your code using markdown format
    at the end of your response.
  • VCoT:
    Question: {question}
    Visualize the state after each reasoning step.
    Please wrap your code using markdown format
    at the end of your response.
Each prompt is designed to elicit specific types of responses while maintaining a consistent structure across different reasoning approaches.
Appendix B Human Annotation
B.1 Dataset Annotation Process
The initial dataset annotation was conducted by authors of computer science Ph.D. candidates with extensive expertise in machine learning and computer vision. We possess comprehensive knowledge of both theoretical foundations and practical applications in the field, ensuring high-quality annotations.
B.2 Main Experiment of Human Evaluation
We carefully constructed a test subset comprising 400 samples, systematically selected from four visual tasks and five different categories (20 samples randomly selected from each combination). Three independent evaluators, all with strong computer science backgrounds but not involved in the initial annotation process, were recruited to assess these samples. These evaluators were specifically chosen for their programming expertise, as the evaluation involved code assessment components.
The evaluation process was structured as follows:
  • Each evaluator independently assessed all 400 samples
  • Results were averaged across all three evaluators to minimize individual bias
B.3 Evaluating LLM’s Scoring Reliability
To assess the reliability of LLMs in scoring tasks, we conducted a evaluation through manual annotation. Our analysis focused on examining whether the scores assigned by LLMs aligned with human judgment and established criteria. The evaluation process followed a structured approach with clearly defined assessment standards.
The following criteria were used to determine the reasonableness of LLM scoring:
  • Consistency: Whether the scores maintained consistency across similar responses
  • Justification: Whether the LLM provided logical explanations for its scoring decisions
  • Alignment: The degree of agreement with human expert scoring
  • Context Understanding: The LLM’s ability to consider contextual factors
We implemented a binary classification system for each scored instance:
  • Reasonable: The score aligns with evaluation criteria
  • Unreasonable: The score shows significant deviation from expected standards
After thorough manual review of the scoring samples, we found that 92% of the LLM’s scoring decisions were deemed reasonable. This high percentage suggests that LLMs demonstrate strong capability in scoring.
B.4 Error Analysis
Our study involved analyzing 100 randomly selected outputs with reasoning processes, split evenly between 50 correct and 50 incorrect cases. The analysis uncovers concerning trends in model performance: of the correct answers, just 33% show both accurate results and valid reasoning, while 17% reach correct conclusions through flawed reasoning - suggesting that correct outputs may not truly indicate effective problem-solving abilities.
We identified four distinct error categories, with visual perception errors being the most significant, comprising 86% of all errors. This notably high percentage indicates a fundamental weakness in how models process and interpret visual information. Instruction comprehension errors and output format violations each make up 6% of the failures, with reasoning process errors accounting for the remaining 2%. Based on this error distribution, it’s clear that substantial improvements to visual perception mechanisms are necessary to achieve reliable visual problem-solving capabilities.
Appendix C Scoring Prompt
Figure 8-15 represent the scoring prompts for different task types and different image types.
Figure 8: Prompt for scoring of multi-modal code generation on bar charts.
Figure 9: Prompt for scoring of multi-modal code generation on line charts.
Figure 10: Prompt for scoring of multi-modal code generation on pie charts.
Figure 11: Prompt for scoring of multi-modal code generation on geometric figures.
Figure 12: Prompt for scoring of multi-modal code generation on function plots.
Figure 13: Prompt for scoring of multi-modal code annotation.
Figure 14: Prompt for scoring of multi-modal code deletion.
Figure 15: Prompt for scoring of multi-modal code modification.

### v2

@@ V-C In-Depth Analysis
Points vs. Lines: Regarding the generation setting for MCE (Annotation) in geometric problems, we further categorize the annotations into two types: angle markings and line labels. As illustrated in Table VI, angle annotations demonstrate notably higher scores compared to line annotations. We attribute this performance difference to the varying complexity of placement requirements: angle markings only need to be positioned near the identified vertex, whereas line labels require to be placed at the midpoint of the line segment. This more stringent positioning requirement for line labels contributes to the increased difficulty in their accurate placement.
V-D Error Analysis
@@
We conduct a manual analysis on 100 randomly selected outputs containing reasoning processes, equally divided between 50 correct and 50 incorrect processes. Our examination reveals concerning patterns in the models’ performance in Fig. 7: among the correct answers, only 33% demonstrates both accurate results and valid reasoning processes, while 17% achieves correct answers through wrong reasoning - indicating that correct outputs do not necessarily reflect sound problem-solving capabilities. We also identify four distinct categories of failures, with visual perception errors emerging as the predominant issue, accounting for 86% of all errors. This remarkably high percentage highlights a crucial limitation in the models’ ability to accurately process and interpret visual elements. Both instruction comprehension errors and output format violations each contribute to 6% of the failures, while reasoning process errors constitute the remaining 2%. This distribution of errors suggests that visual perception mechanisms require substantial improvement for reliable visual problem-solving.
Figure 7: Error analysis of answer and reasoning precess.
Figure 8: Case study for MCG of function plots.
V-E Case Study
As shown in Fig. 8, Qwen2.5-VL-3B-Instruct is presented with an image of a function and asked to provide its mathematical formula in LaTeX format. The correct function, as defined by the reference formula, is $f(x)=x^{2}-1$, which represents a standard parabola shifted one unit down on the y-axis. However, the model under evaluation provides the answer $f(x)=x^{2}$. The evaluation system assigns this response a score of 5 out of 10, classifying it as having the same function type but with a “moderate structural difference”. This case clearly demonstrates the core deficiency identified in our work: a bottleneck in fine-grained visual perception. While the model correctly identifies the general function type as a quadratic parabola ($x^{2}$), showing its capability in high-level visual classification, it fails to perceive the precise details of the graph’s position. Specifically, it does not recognize that the vertex is located at (0, -1) on the coordinate plane, instead defaulting to the most basic representation $f(x)=x^{2}$ with a vertex at the origin. This failure is not one of mathematical reasoning but of visual perception - the model does not correctly identify the vertical shift from the y-axis, as evidenced by the missing 1-1 constant in its output. This case exemplifies how models often grasp the general “what” of an image but struggle with the specific “where” and “how,” which are crucial for accurate mathematical interpretation.
VI related work
@@ V-E Case Study
Multimodal Mathematical Assessment: With the rapid development of MLLMs, numerous benchmarks have emerged to evaluate visual mathematical reasoning capabilities[20, 21, 22, 23, 20, 24, 25, 26, 27, 28]. Previous work [29, 30, 31], such as Inter-gps [9] and GeoQA [32], focuses on geometric problems, while others [33, 34] cover broader tasks including function graphs and statistical charts. Though recent work [5, 7, 8] introduces a “multi-modal input-multi-modal output” paradigm, existing evaluations still assess only final answer accuracy, overlooking the quality of intermediate visual operations during reasoning.
@@
Visual Programming: The decomposition of complex visual tasks into manageable steps through programming interfaces has demonstrated significant potential in advancing visual reasoning capabilities[35, 36, 37, 38]. The core idea behind this approach is to leverage the logical and precise nature of computer programs to transform complex visual problems, which traditionally rely on human intuition and experience, into a series of clear, step-by-step operations. Among them, Visprog [39] and ViperGPT [40] showcase the effectiveness of using MLLMs to generate Python code for sequential visual operations. SKETCHPAD [5] further enables dynamic visual manipulations based on intermediate results, achieving particular success in geometric reasoning through visual operations like drawing auxiliary lines via code. This progress inspires our work to move beyond evaluating just final mathematical solutions and toward assessing the quality and appropriateness of intermediate visual operations in the reasoning process.
VII Future Directions
Enhancing core visual perception and grounding: Our findings show that visual perception is the primary bottleneck, causing the vast majority of errors. Future efforts should be directed towards substantially improving the foundational visual architectures of MLLMs, with a specific emphasis on refining their capabilities for fine-grained feature extraction and precise semantic grounding.
Improving reasoning paradigms for visual tasks: We find that complex prompting strategies like CoT don’t consistently outperform simpler approaches in visual contexts. This suggests a need for reasoning paradigms better suited to visual tasks. Future work could explore strategies that more tightly integrate reasoning steps with visual manipulation, moving beyond purely text-based approaches.
Optimizing visual in-context learning: Our research shows that simply increasing the number of shots doesn’t always improve performance. Future work should investigate more efficient ICL strategies, including methods for selecting the most relevant examples and techniques for compressing visual information to better utilize the model’s context window.
VIII conclusion
@@ VII Future Directions
In this paper, we presented the first fine-grained evaluation benchmark for visual operations in mathematical reasoning tasks. Our framework comprehensively assesses four fundamental visual operations across five types of mathematical visualizations, providing a systematic approach to evaluating MLLMs’ capabilities. Through extensive experiments, we identified significant performance gaps between current models and human, particularly in function plots and MCE (Modification) task. Additionally, we observed that while in-context learning can enhance performance, its effectiveness varies across different tasks and requires careful consideration of shot count. We hope our framework and findings will facilitate the development of more capable models in visual mathematical reasoning.